Racal RA-117
Unit N0115 Log

Serial Number:N0115
Year of Manufacture: 1962 ("15 May 1962" stamped on inside of front panel).
Date of Receipt:2007
State upon Receipt:Non-functional. Some corrosion.

2008 Jun: Dismantled, sub-chassis removed for assessment, cleaning and repairs.
2008 Jun: Stiff KC tuning mechanism. KC tuning mechanism cleaned, bearings freed up with light oil, film reels area cleaned of sticky grease, thrust bearing adjusted.
2008 Jun: VFO switch sloppy action. Shaft removed, collet on shaft cut off due to stuck screw, collet replaced. Hole reemed (manufacturing flaw).
2008 Jul 05: Front panel dismantled, cleaned and polished. Two broken posts on bakelite tuning bezel reglued. New KC and MC windows cut and glued. White position markers on knobs repainted.
2008 Jul 11: Power-Supply/Audio compartment dismantled due to corroded wires and components. Chassis back panel milled out for IEC power connector.
2008 Sep: 3 wide-throat toggle switches removed for use in other unit, replaced with common 1/2"-throat toggles with washers.
2008 Oct 03: Power-Supply/Audio compartment rewiring completed.
  • IEC power connector installed.
  • Power resistors moved to top of chassis.
  • Terminal strips added for filter caps.
  • Tag board added for cathode circuits.
  • C193A (rd-og B+) moved to mode switch area.
  • C204A (gn AVC) moved to terminal strip and ground screw connection added.
  • B+ supply wire connection to harmonics compartment (rd-gn) moved from PS compartment to mode switch.
2008 Oct 17: Recapping of 1st VFO completed. Some resistors replaced also.
2008 Oct 20: 2nd VFO & BFO completed.
2008 Oct 22: VFOs and front panel reinstalled.
2008 Nov 02: IF section recapped.
2008 Nov 06: Mixer sub-chassis completed and reinstalled.
2008 Nov 06: Powered up. Audio OK, BFO OK. 1 MHz, 1.7 MHz present. 2nd VFO tunes 3.48 to 4.31 MHz. 1 MHz harmonics present from 2 to 30 MHz as indicated by another shortwave receiver. No reception.
2008 Nov 06: 2nd VFO adjusted to tune 4.6 MHz at 0 KHz on the scale and 3.6 MHz at 1000 (reverse of what might be expected), by loosening 2 clamp screws on drive shaft and rotating variable capacitor so vanes are meshed at high end of scale.
2008 Nov 07: 1st VFO aligned by pulling V4 and connecting frequency counter to pin 1. Reception is now possible but noise level is very high, sensitivity low.
2008 Nov 08: Working backwards from the audio, shorting stage inputs will quiet the receiver back to the 4th mixer. Following this path, the noise level jumps up around the 1.6 MHz bandpass filter. Mixer sub-chassis removed, some rewiring and more components replaced around the 1.6 MHz filter. Installed C329 is 0.01, not 0.005 as indicated in schematic and parts list. Not much change. Should investigate further.

Unit does seem to work above 1 MHz. Below 1 MHz the noise level is high because the Wideband preselector setting must be used. Tuning the preselector significantly reduces the noise level when it can be used (above 1 MHz).

2008 Nov 11: Most R & C in crystal calibrator replaced. Installed, functions.
2009 Nov: 3-4 mixer subchassis from other unit substituted. Similar performance.
2009 Mar: More R & C in original 3-4 mixer replaced. Substitute 3-4 mixer removed and original re-installed.
2009 Mar: 2nd VFO linearity is out by 6 KHz over the span, particularly at ends of scale (0 & 1000). Improved with some re-positioning of variable cap and vane-tweaking, down to < +/- 1 KHz.
2009 Mar: RE-wiring error discovered: B+ on plate instead of screen of audio output (V22). Corrected, audio much better.
2009 Apr: L302 adjusted for null (1.6 MHz trap into 3rd mixer).
2009 May 31: Used for testing other rebuilt 3/4 mixer subchassis. Upon reassembly, signals are barely audible; traced to open connection of green-mark coax from 2-3 MHz bandpass filter, inside the filter, from repeated torquing. Resoldered.

  .. | Assessment | Unit N16 Log | Unit N0115 Log | Unit N0144 Log | Unit N0714 Log | ..
Racal RA-17 & RA-117